# Safe Flow Matching: Robot Motion Planning with Control Barrier Functions

### v1

@@ 6 Conclusion
In this paper, we tackled the critical challenge of enabling robots to generate safe and efficient trajectories in dynamic and unseen environments, a fundamental requirement for autonomous systems in real-world applications. Existing generative methods, such as diffusion models, while capable of modeling complex trajectory distributions, often incur high computational costs and lack explicit safety guarantees, limiting their practicality in safety-critical scenarios. To address these shortcomings, we proposed SafeFM, a novel motion planning framework that integrates flow matching with CBFs to ensure trajectory safety while maintaining computational efficiency. SafeFM leverages the deterministic and efficient nature of flow matching to transform samples from a simple prior distribution into goal-directed trajectories, embedding safety constraints directly into the learned flow field via CBFs. This approach guarantees that generated trajectories remain within predefined safe regions throughout the planning horizon, even in the presence of unseen obstacles or dynamic constraints. Our experimental evaluations, spanning planar robot navigation, demonstrated SafeFM’s superior performance over conventional flow model planners. Specifically, SafeFM achieved higher success rates, eliminated safety violations entirely after CBF correction, and offered significantly faster inference times, making it highly suitable for real-time robotic applications.
Despite its strengths, SafeFM has limitations that warrant consideration. The current formulation relies on predefined safety constraints, which may be challenging to specify in highly complex or unpredictable settings. Additionally, while effective in the evaluated scenarios, further validation is needed to confirm its scalability to more intricate tasks or real-world conditions with greater variability. Looking ahead, future research could focus on dynamically learning or adapting safety constraints to improve flexibility, integrating SafeFM with complementary frameworks like reinforcement learning for enhanced adaptability, or extending its application to a wider range of robotic domains.

### v2

@@ 6 Conclusion
In this paper, we tackled the critical challenge of enabling robots to generate safe and efficient trajectories in dynamic and unseen obstacles, a fundamental requirement for autonomous systems in real-world applications. Existing generative methods, such as diffusion models, while capable of modeling complex trajectory distributions, often incur high computational costs and lack explicit safety guarantees, limiting their practical applications in safety-critical scenarios. To address these shortcomings, we proposed SafeFlow, a novel motion planning framework that integrates flow matching with FMBFs inspired by CBFs. Our proposed SafeFlow ensures trajectory safety while maintaining computational efficiency. Specifically, SafeFlow leverages the deterministic and efficient nature of flow matching to transform samples from a simple prior distribution into goal-directed trajectories, embedding safety constraints directly into the learned flow field via FMBFs. This approach guarantees that generated trajectories remain within safe regions throughout the planning horizon, even in the presence of unseen obstacles or dynamic constraints. Our experimental evaluations, including planar robot navigation and 7-DoF robot manipulation, demonstrated the superior performance of SafeFlow over SOTA planners. These advantages make it highly suitable for real-time safety-critical applications.